\begin{thebibliography}{9999}
\bibitem{AGU22}
Albert, M. H., Gudmundsson, B., Ulfarsson, H. (2022).
Collatz meets Fibonacci.
\emph{Math. Mag.} 95(2): 130--136.
\href{https://doi.org/10.1080/0025570X.2022.2023307}{doi.org/10.1080/0025570X.2022.2023307}.

\bibitem{Eve77}
Everett, C. J. (1977).
Iteration of the number-theoretic function $f(2n)=n$, $f(2n+1)=3n+2$.
\emph{Adv. Math.} 25(1): 42--45.
\href{https://doi.org/10.1016/0001-8708(77)90087-1}{doi.org/10.1016/0001-8708(77)90087-1}.

\bibitem{Lag85}
Lagarias, J. C. (1985).
The $3x+1$ problem and its generalizations.
\emph{Amer. Math. Monthly} 92(1): 3--23.
\href{https://doi.org/10.2307/2322189}{doi.org/10.2307/2322189}.

\bibitem{Lag10}
Lagarias, J. C. (Ed.). (2010).
\emph{The Ultimate Challenge: The $3x+1$ Problem}.
Providence, RI: American Mathematical Society.

\bibitem{OEIS}
OEIS Foundation Inc. (2026).
The On-Line Encyclopedia of Integer Sequences.
Published electronically at
\href{https://oeis.org}{oeis.org}.

\bibitem{Rey26}
Reyes Jim\'enez, M.-A. (2026).
A Fibonacci theorem for Collatz trajectories via modular graph structure.
Published electronically at
\href{https://arxiv.org/abs/2606.02621}{arXiv:2606.02621}.

\bibitem{Ter76}
Terras, R. (1976).
A stopping time problem on the positive integers.
\emph{Acta Arith.} 30(3): 241--252.
\href{https://doi.org/10.4064/aa-30-3-241-252}{doi.org/10.4064/aa-30-3-241-252}.

\bibitem{Win14}
Winkler, M. (2014).
On the structure and the behaviour of Collatz $3n+1$ sequences - Finite subsequences and the role of the Fibonacci sequence.
Published electronically at
\href{https://arxiv.org/abs/1412.0519v1}{arXiv:1412.0519v1}.
\end{thebibliography}

\documentclass[11pt]{article}

\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}
\usepackage{lmodern}
\usepackage[english]{babel}
\usepackage{amsmath,amssymb,amsthm}
\usepackage{booktabs}
\usepackage{geometry}
\usepackage[hidelinks]{hyperref}
\hypersetup{
  pdftitle={Fibonacci Enumeration of Parity Blocks in Collatz Trajectories},
  pdfauthor={Mike Winkler},
  pdfsubject={Collatz parity blocks and Fibonacci enumeration},
  pdfkeywords={Collatz problem, Fibonacci numbers, parity vectors, binary words, finite automata}
}
\usepackage{url}

\newcommand{\N}{{\mathbb N}}
\newcommand{\Z}{{\mathbb Z}}

\newtheorem{thm}{Theorem}[section]
\newtheorem{theorem}[thm]{Theorem}
\newtheorem{lemma}[thm]{Lemma}
\newtheorem{proposition}[thm]{Proposition}
\newtheorem{corollary}[thm]{Corollary}
\theoremstyle{definition}
\newtheorem{definition}[thm]{Definition}

\newtheorem{remark}[thm]{Remark}

\numberwithin{equation}{section}
\numberwithin{table}{section}
\numberwithin{figure}{section}

\title{Fibonacci Enumeration of Parity Blocks in Collatz Trajectories\\[5mm]}

\author{Mike Winkler\\[5mm]
Ruhr University Bochum\\
Faculty of Mathematics, Germany\\
mike.winkler@ruhr-uni-bochum.de}

\date{August 10, 2026\\[5mm]}

\begin{document}
\maketitle
\begin{abstract}
Let $T(n)=n/2$ for even $n$ and $T(n)=(3n+1)/2$ for odd $n$. For an odd
starting number $s$, let $\ell(s)$ be the first positive index $k$ for which
$T^k(s)\equiv3\pmod4$ and $T^{k-1}(s)$ is even. The use of this single local
cut time is the main structural step: it replaces the two block parameters of
an earlier formulation and converts the endpoint condition into a local
condition on consecutive parity symbols. We prove that, for every $r\geq2$,
the starting numbers $s\equiv1\pmod4$ with $\ell(s)=r$ form exactly
$F_{r-1}$ residue classes modulo $2^{r+2}$, while for every $r\geq3$ the
starting numbers $s\equiv3\pmod4$ with $\ell(s)=r$ form exactly $F_r-1$
such classes. Fixing any odd residue class modulo $12$ leaves these counts
unchanged and yields the two Fibonacci formulas conjectured in a 2014 preprint by the author
as special cases. The proof uses the classical finite parity
correspondence of Terras and Everett. A term is congruent to $3$ modulo $4$
exactly when two consecutive parities are $1$, so the cut conditions reduce
to binary words without the factor $11$. More generally, every finite-state
condition on a parity word gives a residue-class counting sequence governed
by a finite transfer matrix. The present Fibonacci formulas are the two-state
instance of this principle. We also show that the exceptional trajectories
for which no cut occurs reach $1$, and that these exceptions have density
zero in every odd residue class modulo $12$.
\end{abstract}

\par\smallskip
\noindent\textit{Keywords.} Collatz problem, Fibonacci numbers, parity vectors, binary words, residue classes, finite automata, transfer matrices.
\par\smallskip
\noindent\textit{2020 Mathematics Subject Classification.} 11B39 (primary); 11B37, 05A15, 68R15 (secondary).

\section{Introduction}\label{sec:intro}

The Collatz function is defined on the positive integers by
\[
T(n)=
\begin{cases}
 n/2 & \text{if $n$ is even},\\[2pt]
 (3n+1)/2 & \text{if $n$ is odd}.
\end{cases}
\]
For $s\in\N$, let $T^0(s)=s$ and $T^k(s)=T(T^{k-1}(s))$, and write
$C(s)=(T^k(s))_{k\geq0}$ for the trajectory of $s$. The Collatz conjecture
asserts that every trajectory reaches $1$. General references are Lagarias
\cite{Lag85,Lag10}.

This paper proves and strengthens two Fibonacci formulas conjectured in a 2014 preprint by the author
\cite{Win14}, where they were verified
computationally for $2\leq h,t\leq50$. The original formulation selected the
starting classes $9$, $3$, and $7$ modulo $12$ because they arose as block
origins in a global decomposition of Collatz trajectories. It also used two
different block parameters and two different endpoint conventions. The main
structural step in the present paper is to isolate a single local cut time
$\ell$. This removes the global decomposition from the counting problem,
replaces both former parameters by one quantity, and makes the endpoint
condition local in the parity word. Once this recoding is established, the
Fibonacci recurrence follows from the avoidance of one binary factor. Fixing
a residue modulo $3$ then gives the corresponding statements for every odd
residue class modulo $12$, and the formulas of \cite{Win14} become immediate
special cases.

The finite parity correspondence used below belongs to the classical
foundations of the subject and goes back to Terras \cite{Ter76} and Everett
\cite{Eve77}. Fibonacci numbers also occur elsewhere in Collatz combinatorics.
Albert, Gudmundsson, and Ulfarsson \cite{AGU22} studied permutations induced
by finite Collatz trajectories. Their Fibonacci enumeration holds through
length $14$, after which additional permutations occur. More recently, Reyes
Jim\'enez \cite{Rey26} proved that exactly $F_{m+1}$ odd integers in
$\{1,\dots,2^m\}$ avoid the residue class $4\pmod6$ during a prescribed
interval of their trajectories. His proof counts paths in a directed modular
graph and identifies the golden ratio as the spectral radius of the relevant
avoidance subgraph. The objects counted here are different: complete residue
classes determined by a local cut in $[3]_4$, for every value of the block
length. Nevertheless, the two results share a finite-state mechanism. A local
condition on a trajectory is encoded by paths in a finite state space and is
then counted by a transfer matrix. In the present setting the state space
reduces to binary parity data and the only internal forbidden factor is $11$.
Section~\ref{sec:finite-state} formulates this common mechanism for arbitrary
finite-state conditions on parity words.

Let $(F_m)_{m\geq0}$ be the Fibonacci sequence defined by
$F_0=0$, $F_1=1$, and $F_{m+1}=F_m+F_{m-1}$ for $m\geq1$. Our main
results are as follows.

\begin{theorem}\label{thm:intro1}
For every $r\geq2$, the set
\[
\{s\equiv1\pmod4:\ell(s)=r\}
\]
is the union of exactly $F_{r-1}$ residue classes modulo $2^{r+2}$.
\end{theorem}

\begin{theorem}\label{thm:intro3}
For every $r\geq3$, the set
\[
\{s\equiv3\pmod4:\ell(s)=r\}
\]
is the union of exactly $F_r-1$ residue classes modulo $2^{r+2}$.
\end{theorem}

The proof has two ingredients. The classical finite parity correspondence
\cite{Ter76,Eve77} identifies binary words of length $k$ with residue classes
modulo $2^k$. The elementary recoding lemma below identifies terms in
$[3]_4$ with occurrences of the factor $11$ in the parity word. The block
classes are therefore counted by words without adjacent ones. A fixed
residue modulo $3$ is independent of the parity word, so the Chinese
remainder theorem transfers the same count to each odd residue class
modulo $12$.

Throughout the paper, $[a]_m$ denotes the residue class
$\{n\in\N:n\equiv a\pmod m\}$.

\section{Parity vectors and binary words}\label{sec:parity}

For $s\in\N$ and $k\geq0$, put
\[
p_k(s)=T^k(s)\bmod2\in\{0,1\},
\qquad
P_k(s)=(p_0(s),\dots,p_{k-1}(s))\in\{0,1\}^k.
\]
The vector $P_k(s)$ is the parity vector of $s$ of length $k$.

The following finite parity correspondence appears in the early work of
Terras \cite{Ter76} and Everett \cite{Eve77}. We include a proof in the
notation used here.

\begin{lemma}[Terras--Everett]\label{lem:terras}
For every $k\geq1$, the parity vector $P_k(s)$ depends only on the residue
of $s$ modulo $2^k$, and the induced map
\[
\Z/2^k\Z\longrightarrow\{0,1\}^k,
\qquad [s]\longmapsto P_k(s),
\]
is a bijection.
\end{lemma}

\begin{proof}
We use induction on $k$. The assertion is clear for $k=1$. Suppose it holds
for $k$. The map
\[
\Phi:\Z/2^{k+1}\Z\longrightarrow\{0,1\}\times\Z/2^k\Z,
\qquad [s]\longmapsto(p_0(s),[T(s)]),
\]
is well-defined. Indeed, replacing $s$ by $s+2^{k+1}m$ changes $T(s)$ by
$2^km$ in the even case and by $3\cdot2^km$ in the odd case. The map is
injective. If two inputs have the same parity and their images under $T$
are congruent modulo $2^k$, division by $2$ in the even case, or
multiplication by the inverse of $3$ modulo $2^{k+1}$ in the odd case
($3$ is odd and hence invertible modulo $2^{k+1}$), shows that the inputs
are congruent modulo $2^{k+1}$. Since the domain and
codomain have the same cardinality, $\Phi$ is bijective. The induction
hypothesis applied to $T(s)$ completes the proof.
\end{proof}

\begin{lemma}[Recoding lemma]\label{lem:recode}
Let $n$ be odd. Then
\begin{align*}
n\equiv1\pmod4&\iff T(n)\text{ is even},\\
n\equiv3\pmod4&\iff T(n)\text{ is odd}.
\end{align*}
Consequently, for every $s\in\N$ and $k\geq0$,
\begin{align*}
T^k(s)\equiv3\pmod4
&\iff p_k(s)=p_{k+1}(s)=1,\\
T^k(s)\equiv1\pmod4
&\iff p_k(s)=1,\quad p_{k+1}(s)=0.
\end{align*}
\end{lemma}

\begin{proof}
If $n=4x+1$, then $T(n)=6x+2$ is even. If $n=4x+3$, then
$T(n)=6x+5$ is odd.
\end{proof}

Thus a term in $[3]_4$ corresponds to an occurrence of $11$ in the parity
word, while a term in $[1]_4$ corresponds to an occurrence of $10$.

\begin{corollary}\label{cor:allones}
For every $k\geq1$,
\[
p_0(s)=\cdots=p_{k-1}(s)=1
\iff s\equiv-1\pmod{2^k}.
\]
In particular, every positive integer has an even term in its trajectory.
\end{corollary}

\begin{proof}
The first assertion follows by induction. If it holds for $k$, then
$p_0(s)=\cdots=p_k(s)=1$ is equivalent to $s$ being odd and
$T(s)\equiv-1\pmod{2^k}$. Hence
$3s+1\equiv-2\pmod{2^{k+1}}$, which is equivalent to
$s\equiv-1\pmod{2^{k+1}}$. For $2^k>s+1$, the congruence
$s\equiv-1\pmod{2^k}$ is impossible.
\end{proof}

\begin{lemma}\label{lem:words}
For $m\geq0$, the number of binary words of length $m$ without the factor
$11$ is $F_{m+2}$.
\end{lemma}

\begin{proof}
Let $g_m$ denote this number. We have $g_0=1$ and $g_1=2$. A word counted
by $g_m$ ends either in $0$, preceded by a word counted by $g_{m-1}$, or in
$01$, preceded by a word counted by $g_{m-2}$. Thus
$g_m=g_{m-1}+g_{m-2}$, and the stated formula follows.
\end{proof}

\section{Cut times and parity blocks}\label{sec:blocks}

For a starting number in $[3]_4$, the trajectory initially contains a
finite run of consecutive odd terms. We record where this run ends.

\begin{definition}\label{def:a}
For $s\equiv3\pmod4$, define
\[
a(s)=\min\{k\geq1:T^k(s)\equiv1\pmod4\}.
\]
\end{definition}

The value $a(s)$ is finite. Otherwise the parity word would consist only of
ones, contrary to Corollary~\ref{cor:allones}. By
Lemma~\ref{lem:recode}, the initial run of ones in the parity word is
$p_0(s)=\cdots=p_{a(s)}(s)=1$, followed by $p_{a(s)+1}(s)=0$; thus
$a(s)+1$ is the length of the initial run.

\begin{definition}[Cut time]\label{def:ell}
Let $s$ be odd. Define
\[
\ell(s)=\min\{k\geq1:T^k(s)\equiv3\pmod4
\text{ and }T^{k-1}(s)\text{ is even}\}.
\]
If the set is empty, put $\ell(s)=\infty$. When $\ell(s)<\infty$, the
finite sequence
\[
B(s)=(T^0(s),\dots,T^{\ell(s)-1}(s))
\]
is the parity block of $s$. The term $T^{\ell(s)}(s)$ is its cut term.
\end{definition}

\begin{lemma}[First-cut characterization]\label{lem:firstcut}
Let $s$ be odd, with the convention that the minimum of an empty set is
infinity.
\begin{enumerate}
\item If $s\equiv1\pmod4$, then
\[
\ell(s)=\min\{k\geq1:T^k(s)\equiv3\pmod4\}.
\]
\item If $s\equiv3\pmod4$, then
\[
\ell(s)=\min\{k>a(s):T^k(s)\equiv3\pmod4\}.
\]
\end{enumerate}
\end{lemma}

\begin{proof}
If the relevant set on the right is empty, then no cut can occur, so both
sides are infinite. Suppose first that $s\equiv1\pmod4$, and let
$k\geq1$ be the first index for which $T^k(s)\equiv3\pmod4$. If $T^{k-1}(s)$ were odd, then
Lemma~\ref{lem:recode} would give $T^{k-1}(s)\equiv3\pmod4$. For $k=1$
this contradicts $s\equiv1\pmod4$, and for $k>1$ it contradicts the
minimality of $k$. Hence $T^{k-1}(s)$ is even, so $k=\ell(s)$.

Now suppose that $s\equiv3\pmod4$, and let $k>a(s)$ be the first later
index for which $T^k(s)\equiv3\pmod4$. If $T^{k-1}(s)$ were odd, then
Lemma~\ref{lem:recode} would give $T^{k-1}(s)\equiv3\pmod4$. Since
$T^{a(s)}(s)\equiv1\pmod4$, we have $k-1>a(s)$, contrary to the
minimality of $k$. Thus $T^{k-1}(s)$ is even, and again $k=\ell(s)$.
\end{proof}

This convention is uniform: the cut term is not part of the block. By
Lemma~\ref{lem:firstcut}, a finite cut time satisfies $\ell(s)\geq2$ when
$s\equiv1\pmod4$, since $T(s)$ is even. When $s\equiv3\pmod4$, the first
possible cut occurs after $T^{a(s)}(s)\equiv1\pmod4$ and its even
successor, so $\ell(s)\geq a(s)+2\geq3$. The ranges in
Theorems~\ref{thm:intro1} and~\ref{thm:intro3} therefore include every
possible finite value.

For the three starting classes used in \cite{Win14},
Lemma~\ref{lem:firstcut} identifies the old parameters as
$h=\ell(s)$ on $[9]_{12}$ and $t=\ell(s)-1$ on
$[3]_{12}\cup[7]_{12}$.

The cut time $\ell(s)$ need not be finite. Such exceptions are harmless for
the Collatz conjecture.

\begin{lemma}[Exceptional trajectories]\label{lem:exceptional}
If $s$ is odd and $\ell(s)=\infty$, then the trajectory of $s$ reaches $1$.
\end{lemma}

\begin{proof}
By Lemma~\ref{lem:firstcut}, no later term belongs to $[3]_4$ when
$s\equiv1\pmod4$, and no term after the initial odd run belongs to
$[3]_4$ when $s\equiv3\pmod4$. Every subsequent odd term is therefore
congruent to $1$ modulo $4$. If $o>1$ is such a term, the next odd term is
\[
\frac{3o+1}{2^j}
\]
for some $j\geq2$, and hence is at most $(3o+1)/4<o$. The positive odd
terms decrease strictly until they reach $1$.
\end{proof}

\begin{remark}\label{rem:local}
Parity blocks are local orbit segments. Distinct blocks may overlap after
two trajectories merge, and a trajectory may enter a block after its first
term. These facts do not affect the residue-class enumeration below, which
depends only on the parity word from the chosen starting number to its
cut term. A global segment decomposition is therefore unnecessary for
the present paper.
\end{remark}

\section{Fibonacci enumeration}\label{sec:counting}

By Lemma~\ref{lem:recode},
\begin{equation}\label{eq:start}
\begin{gathered}
s\equiv1\pmod4
\iff p_0(s)=1,\quad p_1(s)=0,\\
s\equiv3\pmod4
\iff p_0(s)=p_1(s)=1.
\end{gathered}
\end{equation}

\begin{theorem}\label{thm:count1}
For every $r\geq2$, the set
\[
\mathcal B_1(r)=\{s\equiv1\pmod4:\ell(s)=r\}
\]
is the union of exactly $F_{r-1}$ residue classes modulo $2^{r+2}$.
The classes obtained for different values of $r$ are disjoint.
\end{theorem}

\begin{proof}
By Lemmas~\ref{lem:firstcut} and~\ref{lem:recode}, the condition
$\ell(s)=r$ is equivalent to
\begin{equation}\label{eq:ell1}
(p_k(s),p_{k+1}(s))\neq(1,1)\quad(1\leq k\leq r-1),
\qquad p_r(s)=p_{r+1}(s)=1.
\end{equation}
Together with \eqref{eq:start}, the parity conditions determine a word
$(p_0,\dots,p_{r+1})$ of length $r+2$. By
Lemma~\ref{lem:terras}, each admissible word determines one residue class
modulo $2^{r+2}$.

It remains to count the admissible words. We have $p_0=1$, $p_1=0$, and
$p_r=p_{r+1}=1$. For $r\geq3$, condition \eqref{eq:ell1} forces
$p_{r-1}=0$. For $r=2$, the same forced zero is $p_1$. Thus there is
exactly one word for $r=2$ and for $r=3$, in accordance with
$F_1=F_2=1$. If $r\geq4$, the free letters are
$p_2,\dots,p_{r-2}$, a word of length $r-3$ without the factor $11$.
Lemma~\ref{lem:words} gives $F_{r-1}$ choices. Distinct values of $r$ give
disjoint classes because $\ell$ is a function.
\end{proof}

\begin{theorem}\label{thm:count3}
For every $r\geq3$, the set
\[
\mathcal B_3(r)=\{s\equiv3\pmod4:\ell(s)=r\}
\]
is the union of exactly $F_r-1$ residue classes modulo $2^{r+2}$.
The classes obtained for different values of $r$ are disjoint.
\end{theorem}

\begin{proof}
Put $a=a(s)$. By Lemmas~\ref{lem:firstcut} and~\ref{lem:recode}, the
conditions $a(s)=a$ and $\ell(s)=r$ are equivalent to
\begin{equation}\label{eq:ell3}
\begin{gathered}
p_0(s)=\cdots=p_a(s)=1,\qquad p_{a+1}(s)=0,\\
(p_k(s),p_{k+1}(s))\neq(1,1)
\quad(a+1\leq k\leq r-1),\\
p_r(s)=p_{r+1}(s)=1.
\end{gathered}
\end{equation}
Here $a\geq1$ because $s\equiv3\pmod4$, and
$p_{a+1}=0$ together with $p_r=1$ forces $r\geq a+2$; hence
$1\leq a\leq r-2$. The value of $a$ is the length of the initial run of
ones minus one, so it is determined by the word. By
Lemma~\ref{lem:terras}, every admissible word of length $r+2$ gives one
residue class modulo $2^{r+2}$.

Fix $a$. Since $p_r=1$, condition \eqref{eq:ell3} for $k=r-1$ forces
$p_{r-1}=0$. If $r=a+2$ or $r=a+3$, no free letter remains and there is
exactly one word; note that $F_{r-a-1}=1$ in these cases. If
$r\geq a+4$, the free letters are $p_{a+2},\dots,p_{r-2}$, a word of
length $r-a-3$ without the factor $11$. Their number is $F_{r-a-1}$ by
Lemma~\ref{lem:words}. Summing over $a$ gives
\[
\sum_{a=1}^{r-2}F_{r-a-1}
=
\sum_{j=1}^{r-2}F_j
=F_r-1.
\]
Disjointness follows from the definition of $\ell$.
\end{proof}

\begin{corollary}[Refinement modulo $12$]\label{cor:mod12}
Let $c$ be an odd residue class modulo $12$.
\begin{enumerate}
\item If $c\equiv1\pmod4$, then for every $r\geq2$ the set
\[
\{s\equiv c\pmod{12}:\ell(s)=r\}
\]
is the union of exactly $F_{r-1}$ residue classes modulo
$12\cdot2^r$.
\item If $c\equiv3\pmod4$, then for every $r\geq3$ the same set is the
union of exactly $F_r-1$ residue classes modulo $12\cdot2^r$.
\end{enumerate}
\end{corollary}

\begin{proof}
A class modulo $2^{r+2}$ occurring in Theorem~\ref{thm:count1} or
Theorem~\ref{thm:count3} already fixes the residue modulo $4$. Fixing
$s\equiv c\pmod{12}$ additionally fixes one residue modulo $3$. Since
$3$ and $2^{r+2}$ are coprime, the Chinese remainder theorem refines each
such class to exactly one residue class modulo
$3\cdot2^{r+2}=12\cdot2^r$. The number of classes is unchanged.
\end{proof}

\begin{corollary}[The conjectures of \cite{Win14}]\label{cor:old}
For every $r\geq2$, the starting numbers $s\equiv9\pmod{12}$ with
$\ell(s)=r$ form exactly $F_{r-1}$ residue classes modulo
$12\cdot2^r$. For every $r\geq3$, the starting numbers
$s\equiv3,7\pmod{12}$ with $\ell(s)=r$ form exactly $2F_r-2$ residue
classes modulo the same modulus.
\end{corollary}

\begin{proof}
The first assertion is Corollary~\ref{cor:mod12} with $c=9$. The second is
the disjoint union of its cases $c=3$ and $c=7$.
\end{proof}

Table~\ref{tab:classes} lists the first classes from
Corollary~\ref{cor:old}. Under the notation of \cite{Win14}, the first half
is indexed by $h=r$ and the second by $t=r-1$. The first sequence satisfies
$F_{r-1}=\mathrm{A000045}(r-1)$ for $r\geq2$, while
$2F_r-2=\mathrm{A019274}(r)$ for $r\geq3$ in the OEIS \cite{OEIS}.

\begin{table}[ht]
\caption{The first finite-block classes from Corollary~\ref{cor:old}; the
modulus is shown in parentheses.}
\label{tab:classes}
\centering
\begin{tabular}{clc}
\toprule
$r$ & classes in $[9]_{12}$ modulo $12\cdot2^r$ & $F_{r-1}$\\
\midrule
2 & $9\ (48)$ & 1\\
3 & $93\ (96)$ & 1\\
4 & $33,\ 165\ (192)$ & 2\\
5 & $81,\ 117,\ 237\ (384)$ & 3\\
6 & $129,\ 333,\ 405,\ 561,\ 645\ (768)$ & 5\\
\midrule
$r$ & classes in $[3]_{12}\cup[7]_{12}$ modulo $12\cdot2^r$
& $2F_r-2$\\
\midrule
3 & $27,\ 91\ (96)$ & 2\\
4 & $19,\ 39,\ 103,\ 147\ (192)$ & 4\\
5 & $55,\ 67,\ 111,\ 183,\ 195,\ 235,\ 363,\ 367\ (384)$ & 8\\
\bottomrule
\end{tabular}
\end{table}

\begin{corollary}\label{cor:density}
We have
\[
\sum_{r\geq2}\frac{F_{r-1}}{2^r}=1,
\qquad
\sum_{r\geq3}\frac{F_r-1}{2^r}=1.
\]
Consequently, for every odd residue class $c$ modulo $12$, the
finite-block classes have full relative density in $[c]_{12}$; equivalently,
their complement in $[c]_{12}$ has natural density zero. Every starting
number in an exceptional set reaches $1$.
\end{corollary}

\begin{proof}
The generating function
\[
\sum_{m\geq1}F_mx^m=\frac{x}{1-x-x^2}
\]
gives $\sum_{m\geq1}F_m2^{-m}=2$. Hence
\[
\sum_{r\geq2}\frac{F_{r-1}}{2^r}=1
\]
and
\[
\sum_{r\geq3}\frac{F_r-1}{2^r}
=
\sum_{r\geq3}\frac{F_r}{2^r}
-
\sum_{r\geq3}\frac{1}{2^r}
=
\left(2-\frac12-\frac14\right)-\frac14=1.
\]
Fix an odd residue class $c$ modulo $12$. The class $[c]_{12}$ has density
$1/12$, while each class modulo $12\cdot2^r$ has density
$(12\cdot2^r)^{-1}$. If $c\equiv1\pmod4$, then for fixed $R$ the complement
in $[c]_{12}$ of the classes with $2\leq\ell(s)\leq R$ has density
\[
\frac1{12}\left(1-\sum_{r=2}^{R}\frac{F_{r-1}}{2^r}\right),
\]
which tends to zero. If $c\equiv3\pmod4$, the analogous density is
\[
\frac1{12}\left(1-\sum_{r=3}^{R}\frac{F_r-1}{2^r}\right),
\]
which also tends to zero. The exceptional set is contained in every finite
remainder and therefore has upper density zero. The final assertion is
Lemma~\ref{lem:exceptional}.
\end{proof}

\section{Finite-state formulation}\label{sec:finite-state}

The preceding proofs use only one forbidden factor, but the underlying
mechanism is more general. The parity correspondence converts any condition
on the first $k$ parity symbols into a condition on residue classes modulo
$2^k$. If the word condition can be recognized with finitely many states, its
count is therefore controlled by a finite transfer matrix.

\begin{proposition}[Finite-state parity principle]\label{prop:finite-state}
Let $\mathcal A$ be a finite deterministic partial automaton over the alphabet
$\{0,1\}$, with a specified initial state and a specified set of accepting
states. Let $a_k$ be the number of binary words of length $k$ accepted by
$\mathcal A$. Then exactly $a_k$ residue classes modulo $2^k$ have a parity
vector of length $k$ accepted by $\mathcal A$.

If $M$ is the transfer matrix of $\mathcal A$, whose $(i,j)$ entry is the
number of symbols taking state $i$ to state $j$, then
\[
 a_k=u^{\mathsf T}M^k v,
\]
where $u$ is the indicator vector of the initial state and $v$ is the
indicator vector of the accepting states. Consequently, $(a_k)$ satisfies a
linear recurrence with constant coefficients determined by the
characteristic polynomial of $M$.
\end{proposition}

\begin{proof}
By Lemma~\ref{lem:terras}, the map from residue classes modulo $2^k$ to binary
parity words of length $k$ is a bijection. Hence the number of residue classes
whose parity words are accepted by $\mathcal A$ is exactly the number $a_k$ of
accepted words.

Because the automaton is deterministic, every word determines at most one
path from the initial state. The matrix product $u^{\mathsf T}M^k v$ therefore
counts exactly the accepted paths of length $k$, and hence the accepted words.
If
\[
 \chi_M(x)=x^d+c_{d-1}x^{d-1}+\cdots+c_0
\]
is the characteristic polynomial, the Cayley--Hamilton theorem gives
\[
 M^{k+d}+c_{d-1}M^{k+d-1}+\cdots+c_0M^k=0.
\]
Multiplication by $u^{\mathsf T}$ and $v$ yields the corresponding recurrence
for $a_k$.
\end{proof}

For the language of binary words avoiding $11$, two states suffice: one
records that the current word ends in $0$ (or is empty), and the other that it
ends in $1$. The transfer matrix is
\[
 M=\begin{pmatrix}1&1\\1&0\end{pmatrix},
\]
with characteristic polynomial $x^2-x-1$. Thus Proposition~\ref{prop:finite-state}
recovers the Fibonacci recurrence in Lemma~\ref{lem:words}; the Perron root is
the golden ratio $\varphi=(1+\sqrt5)/2$. Theorems~\ref{thm:count1} and
\ref{thm:count3} add fixed prefix and suffix conditions and, in the second
case, a summation over the initial run length, but their free words are
counted by this same two-state system.

This also clarifies the relation to the modular-graph method of Reyes
Jim\'enez \cite{Rey26}. His states record residues modulo $6$, whereas the
states above record local parity information. Thus his theorem is not a
direct corollary of Proposition~\ref{prop:finite-state}. Both constructions,
however, turn a local orbit restriction into path counting in a finite state
space and then into a transfer-matrix problem. This is the common structural
source of the Fibonacci behavior.

\section{Concluding remarks}\label{sec:conclusion}

Theorems~\ref{thm:count1} and~\ref{thm:count3} give more than numerical
identities. They give bijective descriptions of the block classes in terms of
binary words without adjacent ones. In Theorem~\ref{thm:count1}, the free
word has length $r-3$. In Theorem~\ref{thm:count3}, one first chooses the
length of the initial run and then a free word without adjacent ones. The
Fibonacci recurrence is therefore intrinsic to the local block structure.
The role of the new cut time $\ell$ is essential here: it is the device that
turns the former block conditions into one uniform finite-state condition.
Fixing a residue modulo $3$ merely refines each class by the Chinese remainder
theorem and does not alter the word count.

Proposition~\ref{prop:finite-state} places this enumeration in a wider but still
elementary framework. The present paper is the two-state case in which one
forbidden factor yields the Fibonacci matrix. More complicated local orbit
conditions can be represented by larger automata and therefore lead to other
linear recurrences. A natural next problem is to determine which modular or
orbit-theoretic conditions give particularly small automata and hence
explicit counting formulas.

A second open problem is arithmetic rather than combinatorial. Words without
adjacent ones are equivalent to subsets with no consecutive elements and to
Zeckendorf-type data. It would be useful to express the corresponding residue
representatives modulo $2^{r+2}$ directly in such terms and then to describe
their refinements modulo $12\cdot2^r$.

Finally, the same finite parity mechanism is available for related shortcut
maps with an odd multiplier, since odd multipliers remain invertible modulo
powers of $2$. Whether the cut-time construction developed here admits a
comparably natural and useful extension to maps of the form
$n\mapsto(qn+1)/2$ is left for separate investigation. No claim about the
global Collatz conjecture follows from the finite-state enumeration.

\section*{Declaration on the use of AI tools}
Generative AI tools were used as auxiliary tools during the development of
this work for exploratory mathematical discussion, consistency checks, and
editorial revision. The author independently verified all mathematical
statements, proofs, computations, and references and assumes full
responsibility for the content.

\end{document}